solve for 'x' :
4logx-logx^3 +log2 = log(6-x)

Answer by robertb(5830)

log(x^4) - log(x^3)+log2 = log(6-x).
log(x^4/x^3)+log2 = log(6-x),
logx+log2 = log(6-x),
log(2x) = log(6-x),
2x = 6-x, since log is a 1-to-1 function. Hence
3x = 6,
x = 2. This satisfies the original equation, and thus is the final answer.